# A faster algorithm for Vertex Cover parameterized by solution size

### v6

@@ 10 Branching on 6-vertices
Simplify $G$ If $\operatorname{maxdeg}(G)\leq 5$, then run the cheaper of Branch5-3 or the MaxIS-5 algorithm. Apply Section 5 to a vertex of degree at least $6$
Algorithm 8 Function Branch6-1$(G,k)$
@@
Given the potential branch-seqs from Section 5, we have the following constraints:
@@
1 Apply miscellaneous branching rules to ensure $G$ has nice structure e.g. max degree 6. Split on a 6-vertex $u$ where $G-N[u]$ is 2-vul. Branch on vertices $u,v$ with $\deg_{G}(u)=6,\deg_{G-u}(v)\geq 5$, and either $R(u,v)\geq 1$ or $\operatorname{shad}(u,N[v])\leq 0$. Branch on non-adjacent 6-vertices $x_{1},x_{2},x_{3}$ with $R(x_{1},x_{2},x_{3})\geq 1$. Branch on a 6-vertex $u$ with six 6-neighbors, or five 6-neighbors and one 5-neighbor. if $L=2$ then
2        Branch on non-adjacent 6-vertices $x_{1},x_{2}$ with $R(x_{1},N[x_{2}])\geq 1$. Branch on a 5-vertex $u$ which has two 6-neighbors and three 5-neighbors,            UNLESS each of these 6-neighbors itself has at least four 5-neighbors.
Run the cheaper of Branch6-$(L-1)(G,k)$ or the MaxIS algorithm
Algorithm 9 Function Branch6-$L$$(G,k)$ for $L=2,3$
@@ 10.1 Regularizing the graph and structural properties
There is a grab bag of isolated local obstructions that we will remove. We say that $G$ is in 6-base position (or 6BP for short) if the following conditions hold:
@@
  • Every 6-vertex $u$ has $\operatorname{shad}(N[u])\geq 2$ and $R(u)=0$.
  • Every 5-vertex $u$ has $\operatorname{shad}(N[u])\geq 1$ and $R(u)=0$.
  • There are no vertices $u,v$ with $\operatorname{codeg}(u,v)\geq 2$ and $R(u,v)+R(N(u)\cap N(v))+\operatorname{codeg}(u,v)\geq 5$.
Observation 61.
  • If $G$ is in 5BP, then any 6-vertex $u$ and vertex $v\not\sim u$ have $\operatorname{codeg}(u,v)\leq\deg(v)-3$.
  • If $G$ is in 5BP, then any 5-vertex $u$ and vertex $v\not\sim u$ have $\operatorname{codeg}(u,v)\leq 2$.
@@
In the first case, if $\operatorname{codeg}(u,v)\geq\deg(v)-2$, then $\operatorname{shad}(N[u])\leq 1$. In the second case, we may suppose that also $\deg(v)\leq 5$ (else it is covered by the first case); then the analysis is the same as Proposition 47. In the third case, if $\operatorname{shad}(u,v)\leq 1$ or $\operatorname{shad}(u,N[v])\leq 0$, we would branch at Line 3. In the final case, we have $\operatorname{shad}(x_{1},x_{2},x_{3})\geq\operatorname{shad}(x_{1},x_{2})-1\geq 1$. If $\operatorname{shad}(x_{1},x_{2},x_{3})\leq 1$, then we would branch at Line 4. ∎
@@
Suppose $u$ has 6-neighbors $v_{1},v_{2},v_{3}$. Let $v_{4},v_{5}$ be the other two neighbors of $u$. They must have degree at least 4, else $R(u)\geq 1$ contradicting that $G$ is in 6BP. If $v_{1},v_{2},v_{3}$ are independent, then $R(v_{1},v_{2},v_{3})\geq 1$ due to 2-vertex $u$, and we would apply Line 4 to $v_{1},v_{2},v_{3}$. So we may suppose without loss of generality that $v_{1}\sim v_{2}$.
Now suppose for contradiction that $\deg(v_{4})\geq 5$. If $v_{3}\sim v_{4}$, then $R(v_{1},v_{2})\geq 1$ due to 3-triangle $u,v_{3},v_{4}$, and we would branch on $v_{1},v_{2}$ at Line 3. Otherwise, if $v_{3}\not\sim v_{4}$, then $R(v_{3},v_{4})\geq 1$ due to 3-triangle $u,v_{1},v_{2}$, and we would branch on $v_{3},v_{4}$ at Line 3. ∎
@@
For $i\geq 0$, let $A_{6i}$ be the set of 6-vertices with exactly $i$ 5-neighbors and let $A_{5i}$ be the set of 5-vertices with exactly $i$ 6-neighbors. By Proposition 62, $A_{54}=A_{55}=\emptyset$ and any vertex in $A_{53}$ has at least two 4-neighbors. Also let $A_{52}^{\prime}\subseteq A_{52}$ be the set of 5-vertices with two 6-neighbors and at least one 4-neighbor. By double-counting edges between $V_{5}$ and $V_{6}$, we have:
@@
Now consider the edges between $V_{4}$ and $V_{5}\cup V_{6}$. Each vertex in $A_{60}\cup A_{61}$ contributes one such edge, each vertex in $A_{52}^{\prime}$ contributes at least one edge, each vertex in $A_{53}$ contributes at least two edges by Proposition 62, and each vertex in $V_{4}$ has at most four such edges. So we get:
@@ 10.2 Basic analysis
Line 1. We first simplify $G$. If a vertex $u$ of degree 5 or 6 has $\operatorname{shad}(N[u])\leq 5-\deg(u)$, we apply Lemma 26. If a vertex $u$ of degree 5 or 6 has $R(u)\geq 1$, we split on $u$ with branch-seq $[(0.5,2),(2,5)]$. If a vertex has degree at least 7, we apply Section 5. If there are vertices $u,v$ with $R(u,v)+R(N(u)\cap N(v))+\operatorname{codeg}(u,v)\geq 5$, we apply (A3).
@@
Line 2. Since $G$ is in 6BP, subproblems $G-u$ and $G-N[u]$ have drop $(0.5,1)$ and $(2.5,6)$ directly, and we can apply Proposition 31 to the latter problem. We need:
@@
Line 3. Split on $u$, and in subproblem $G-u$, either apply Lemma 26 or split on $v$. Since $G$ is in 6BP, subproblems $G-N[u],G-u$ have drops $(2.5,6),(0.5,1)$ directly and are simplified. We need:
@@
Line 4. Split on $x_{1},x_{2},x_{3}$, getting subproblems $G_{1}=G-N[x_{1}],G_{2}=G-x_{1}-N[x_{2}],G_{3}=G-\{x_{1},x_{2}\}-N[x_{3}],G_{4}=G-\{x_{1},x_{2},x_{3}\}$. By Observation 61 we have $\operatorname{minsurp}(G_{1})\geq 2,\operatorname{minsurp}(G_{2})\geq 1,\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(N[x_{3}])-2\geq 0$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(x_{1},x_{2})-1\geq 1$. Because $S(G_{4})\geq 1$, these subproblems $G_{1},G_{2},G_{3},G_{4}$ have drops $(2.5,6),(3,7),(3.5,8),(2,5)$. We need:
@@
Line 7. Since $G$ is in 6BP, subproblems $G-N[x_{1}]$ and $G-x_{1}-N[x_{2}]$ and $G-\{x_{1},x_{2}\}$ have drops $(2.5,6),(3,7),(1,2)$ directly, and $G-x_{1}-N[x_{2}]$ has drop $(3,8)$ after simplification. We need:
@@ 10.3 Line 5: branching on a 6-vertex with high-degree neighbors
First observe that $H$ cannot have any triangle $v_{1},v_{2},v_{3}$; for, in that case, we would have $R(v_{4},v_{5})\geq 1$ where $v_{4},v_{5}$ are any other 6-neighbors of $u$; we would branch on $v_{4},v_{5}$ at Line 3.
@@
If $H$ is a matching with exactly one edge $(v_{1},v_{2})$, with three other 6-neighbors $v_{3},v_{4},v_{5}$, we would have $R(v_{3},v_{4},v_{5})\geq 1$ due to the 3-triangle $u,v_{1},v_{2}$ and we would branch Line 4.
If $H$ is a matching with exactly two edges, suppose without loss of generality they are $(v_{1},v_{2}),(v_{3},v_{4})$ where $\deg(v_{4})=6$. If $\deg(v_{5})=\deg(v_{6})=6$, then again have $R(v_{4},v_{5},v_{6})\geq 1$ due to 3-triangle $u,v_{1},v_{2}$, and we would branch at Line 4. So suppose without loss of generality that $\deg(v_{5})=5$, in which case the ordering $v_{1},\dots,v_{6}$ satisfies Case II.
@@
Next suppose $H$ is not a matching. First, suppose some $v_{4}\in H$ has neighbors $v_{5},v_{6}\in H$ where one of the vertices $v_{4},v_{5},v_{6}$ has degree 5. Since $H$ has no triangle, we order the other vertices as $v_{1},v_{2},v_{3}$ where $v_{2}\not\sim v_{3}$, satisfying Case III.
Finally, suppose $v_{4}\in H$ has neighbors $v_{5},v_{6}\in H$ where $\deg(v_{4})=\deg(v_{5})=\deg(v_{6})=6$. The induced subgraph $H^{\prime}=G[\{v_{1},v_{2},v_{3}\}]$ must be a matching, as otherwise it would be covered by the argument in the preceding paragraph. So let $x$ be a vertex which is isolated in $H^{\prime}$. If $\deg(x)=6$, we set $v_{1}=x$ and $v_{2}$ to be the other 5-vertex (if any) and $v_{3}$ to be the 6-vertex. If $\deg(x)=5$, we set $v_{1},v_{2}$ to be the 6-neighbors of $u$ and $v_{3}=x$. ∎
@@
Case Ia: $\boldsymbol{G-\{v_{1},v_{2}\}-N[v_{3}]}$ is 3-vul. Split on $v_{1},\dots,v_{4}$, generating subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}],G_{4}=G-\{v_{1},v_{2},v_{3}\}-N[v_{4}],G_{5}=G-\{v_{1},v_{2},v_{3},v_{4}\}$. By Observation 61, we have $\operatorname{minsurp}(G_{1})\geq 2$, $\operatorname{minsurp}(G_{2})\geq 1$, $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(N[v_{3}])-2\geq 0$, $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(N[v_{4}])-3\geq-1$ and $\operatorname{minsurp}(G_{5})\geq\operatorname{shad}(v_{1},v_{2},v_{3})-1\geq 1$. So $G_{1},G_{2},G_{3},G_{4},G_{5}$ have drop $(2.5,6),(3,7),(3.5,8),(3.5,9),(2,4)$ directly. We can apply Proposition 31 to $G_{3}$, and have $S(G_{5})\geq 1$ due to 2-vertex $u$. We need:
@@
Case Ib: Split on $v_{1},v_{2},v_{3},v_{4}$, and apply (A1) to vertex $u$ in the subproblem $G-\{v_{1},v_{2},v_{3},v_{4}\}$. This generates subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}],G_{4}=G-\{v_{1},v_{2},v_{3}\}-N[v_{4}],G_{0}=G-N[u]$ and $G_{5}=G-\{v_{1},v_{2},v_{3},v_{4}\}-N[v_{5},v_{6}]$.
Since $G$ is 6BP, $G_{0},G_{1},G_{2},G_{3}$ have drops $(2.5,6),(2.5,6),(3,7),(3.5,8)$ directly. We must have $\operatorname{minsurp}(G_{4})\geq 0$, as otherwise $\operatorname{shad}(v_{1},v_{2},N[v_{4}])\leq 0$, which would put us into Case Ia (with swapping the order of $v_{3}/v_{4}$). So $G_{4}$ has drop $(4,9)$.
For subproblem $G_{5}$, let $d=\deg_{G-N[v_{5}]}(v_{6})$; by Observation 61 we have $d\geq 3$. We claim that $\operatorname{minsurp}(G_{5})\geq 2-d$. For, consider an indset $I$ with $\operatorname{surp}_{G_{5}}(I)\leq 1-d\leq-2$; necessarily $|I|\geq 2$. Then $\operatorname{surp}_{G-\{v_{1},v_{2}\}-N[v_{5}]}(I\cup\{v_{6}\})\leq(1-d)+(d-1)+1=2$. So $G-\{v_{1},v_{2}\}-N[v_{5}]$ is 3-vul, putting us into Case Ia with alternate vertex ordering $v_{1}^{\prime}=v_{1},v_{2}^{\prime}=v_{2},v_{3}^{\prime}=v_{5}$. So $G_{5}$ is principal with excess two and $\Delta k=10+d\geq 13$ and $\Delta\mu\geq 5$.
@@
Case IIa: $\boldsymbol{G-v_{6}-N[v_{1}]}$ is 3-vul. Split on $v_{6},v_{1},v_{5}$, obtaining subproblems $G_{1}=G-N[v_{6}],G_{2}=G-v_{6}-N[v_{1}],G_{3}=G-\{v_{1},v_{6}\}-N[v_{5}],G_{4}=G-\{v_{1},v_{6},v_{5}\}$. By Observation 61, we have $\operatorname{minsurp}(G_{1})\geq 2,\operatorname{minsurp}(G_{2})\geq 1,\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(v_{6},N[v_{5}])-1\geq 0,\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(v_{1},v_{6})-1\geq 1$. So $G_{1},G_{2},G_{3},G_{4}$ have drops $(2.5,6),(3,7),(3,7),(1.5,3)$ directly. We apply Proposition 31 to $G_{3}$ and we have $S(G_{4})\geq 1$ (due to 3-triangle $u,v_{3},v_{4}$), so overall we need:
@@
Case IIb: Split on $v_{6},v_{4},v_{3}$ and apply (A2) to vertex $v_{1}$ in subproblem $G-\{v_{6},v_{4},v_{3}\}$. Overall, this generates subproblems $G_{1}=G-N[v_{6}],G_{2}=G-v_{6}-N[v_{4}],G_{3}=G-v_{6}-N[v_{3}],G_{4}=G-\{v_{3},v_{4},v_{6},v_{1}\}$ and $G_{5}=G-\{v_{3},v_{4},v_{6}\}-N[v_{1},v_{5}]$. By Observation 61, we have $\operatorname{minsurp}(G_{1})\geq 2$ and $\operatorname{minsurp}(G_{2})\geq 1$ and $\operatorname{minsurp}(G_{3})\geq 1$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(v_{4},v_{6})-2\geq 0$. So the subproblems have drops $(2.5,6),(3,7),(3,7),(2,4)$ directly; subproblem $G_{4}$ has a 2-vertex $u$, so it has drop $(2,5)$ after simplification.
For subproblem $G_{5}$, let $d=\deg_{G-v_{6}-N[v_{1}]}(v_{5})$. We have $d\geq 3$, as otherwise $R(v_{6},N[v_{1}])\geq 1$ and it would be covered in Case IIa. We must have $\operatorname{minsurp}(G_{5})\geq 2-d$; for, if we have an indset $I$ with $\operatorname{surp}_{G_{5}}(I)\leq 1-d\leq-2$, where necessarily $|I|\geq 2$, then $\operatorname{surp}_{G-v_{6}-N[v_{1}]}(I\cup\{v_{5}\})\leq 2$ and $G-v_{6}-N[v_{1}]$ is 3-vul; it would be covered in Case IIa.
@@
Case III. Split on $v_{1},v_{2},v_{3}$, obtaining subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}],G_{4}=G-\{v_{1},v_{2},v_{3}\}$. By Observation 61, we have $\operatorname{minsurp}(G_{1})\geq 2$ and $\operatorname{minsurp}(G_{2})\geq 1$ and $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(v_{1},N[v_{3}])-1\geq 0$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(v_{1},v_{2})-1\geq 1$. So the subproblems have drops $(2.5,6),(2.5,6),(3,7),(1.5,3)$ directly; furthermore, $G_{4}$ has a kite $u,v_{4},v_{5},v_{6}$ and so it has net drop $(2,5)$ by Lemma 9. We need:

### v7

@@ 8 Branching on 6-vertices
1Simplify $G$
2If $\operatorname{maxdeg}(G)\leq 5$, then run the cheaper of Branch5-3 or the MaxIS-5 algorithm.
3Apply Figure 6 to a vertex of degree at least 66
Algorithm 8 Function Branch6-1(G,k)(G,k)
@@
Given the potential branch-seqs from Figure 6, we have the following constraints:
@@
1
2Apply miscellaneous branching rules to ensure $G$ has nice structure.
3Split on a 6-vertex $u$ where GN[u]G-N[u] is 2-vul.
4Branch on vertices $u,v$ with $\deg_{G}(u)=6,\deg_{G-u}(v)\geq 5$ and $S(u,v)+S(u,N[v])\geq 1$.
5Branch on non-adjacent 6-vertices $x_{1},x_{2},x_{3}$ with $S(x_{1},x_{2},x_{3})\geq 1$.
6Branch on a 6-vertex $u$ with six 6-neighbors, or five 6-neighbors and one 5-neighbor.
7if $L=2$ then
8
9  Branch on non-adjacent 6-vertices $x_{1},x_{2}$ with $S(x_{1},N[x_{2}])\geq 1$.
10  Branch on a 5-vertex $u$ which has two 6-neighbors and three 5-neighbors, UNLESS each of these 6-neighbors itself has at least four 5-neighbors.
11
12Run the cheaper of Branch6-(L1)(G,k)(L-1)(G,k) or the MaxIS algorithm
Algorithm 9 Function Branch6-LL(G,k)(G,k) for $L=2,3$
@@ 8.1 Regularizing the graph and structural properties
In Line 1, we will ensure that the following conditions hold:
@@
  • Every 6-vertex $u$ has $\operatorname{shad}(N[u])\geq 2$ and $S(u)=0$.
  • Every 5-vertex $u$ has $\operatorname{shad}(N[u])\geq 1$ and $S(u)=0$.
  • There are no vertices $u,v$ with $\operatorname{codeg}(u,v)\geq 2$ and $S(u,v)+S(N(u)\cap N(v))+\operatorname{codeg}(u,v)\geq 5$.
Observation 62.
  • After Line 1, any 6-vertex $u$ and vertex $v\not\sim u$ have $\operatorname{codeg}(u,v)\leq\deg(v)-3$.
  • After Line 1, any 5-vertex $u$ and vertex $v\not\sim u$ have $\operatorname{codeg}(u,v)\leq 2$.
@@
In the first case, if $\operatorname{codeg}(u,v)\geq\deg(v)-2$, then $\operatorname{shad}(N[u])\leq 1$. In the second case, we may suppose that also $\deg(v)\leq 5$ (else it is covered by the first case); then the analysis is the same as Proposition 48. In the third case, if $\operatorname{shad}(u,v)\leq 1$ or $\operatorname{shad}(u,N[v])\leq 0$, then we would have $S(u,v)\geq 1$ or $S(u,N[v])\geq 1$ (note that $\operatorname{shad}(u,N[v])\geq\operatorname{shad}(N[v])-1\geq 0$), and we would branch at Line 3. In the final case, we have $\operatorname{shad}(x_{1},x_{2},x_{3})\geq\operatorname{shad}(x_{1},x_{2})-1\geq 1$. If $\operatorname{shad}(x_{1},x_{2},x_{3})\leq 1$, then we would branch at Line 4. ∎
@@
Suppose $u$ has 6-neighbors $v_{1},v_{2},v_{3}$. Let $v_{4},v_{5}$ be the other two neighbors of $u$. They must have degree at least 4, else $S(u)\geq 1$ and we would have branched at Line 1. If $v_{1},v_{2},v_{3}$ are independent, then $S(v_{1},v_{2},v_{3})\geq 1$ due to 2-vertex $u$, and we would apply Line 4 to $v_{1},v_{2},v_{3}$. So we may suppose without loss of generality that $v_{1}\sim v_{2}$.
Now suppose for contradiction that $\deg(v_{4})\geq 5$. If $v_{3}\sim v_{4}$, then $S(v_{1},v_{2})\geq 1$ due to 3-triangle $u,v_{3},v_{4}$, and we would branch on $v_{1},v_{2}$ at Line 3. Otherwise, if $v_{3}\not\sim v_{4}$, then $S(v_{3},v_{4})\geq 1$ due to 3-triangle $u,v_{1},v_{2}$, and we would branch on $v_{3},v_{4}$ at Line 3. ∎
@@
For $i\geq 0$, let $A_{6i}$ be the set of 6-vertices with exactly $i$ 5-neighbors and let $A_{5i}$ be the set of 5-vertices with exactly $i$ 6-neighbors. By Proposition 63, $A_{54}=A_{55}=\emptyset$ and any vertex in $A_{53}$ has at least two 4-neighbors. Also let $A_{52}^{\prime}\subseteq A_{52}$ be the set of 5-vertices with two 6-neighbors and at least one 4-neighbor. By counting edges between $V_{5}$ and $V_{6}$, we have:
@@
Now consider the edges between $V_{4}$ and $V_{5}\cup V_{6}$. Each vertex in $A_{60}\cup A_{61}$ contributes one such edge, each vertex in $A_{52}^{\prime}$ contributes at least one edge, each vertex in $A_{53}$ contributes at least two edges, and each vertex in $V_{4}$ has at most four such edges. So we get:
@@ 8.2 Basic analysis
Line 1. We first simplify $G$. If a vertex $u$ of degree 5 or 6 has $\operatorname{shad}(N[u])\leq 5-\deg(u)$, we apply Lemma 26. If a vertex $u$ of degree 5 or 6 has $S(u)\geq 1$, we split on $u$ with branch-seq $[(0.5,2),(2,5)]$. If a vertex has degree at least 7, we apply Figure 6. If there are vertices $u,v$ with $S(u,v)+S(N(u)\cap N(v))+\operatorname{codeg}(u,v)\geq 5$, we apply (B-Shared).
@@
Line 2. After Line 1, subproblems GuG-u and GN[u]G-N[u] have drop $(0.5,1)$ and $(2.5,6)$ directly, and we can apply Proposition 31 to the latter problem. We need:
@@
Line 3. Split on $u$, and in subproblem GuG-u, either apply Lemma 26 or split on $v$. After Line 1, subproblems GN[u],GuG-N[u],G-u have drops $(2.5,6),(0.5,1)$ directly and are simplified. We need:
@@
Line 4. Split on $x_{1},x_{2},x_{3}$, getting subproblems $G_{1}=G-N[x_{1}],G_{2}=G-x_{1}-N[x_{2}],G_{3}=G-\{x_{1},x_{2}\}-N[x_{3}],G_{4}=G-\{x_{1},x_{2},x_{3}\}$. By Observation 62 we have $\operatorname{minsurp}(G_{1})\geq 2,\operatorname{minsurp}(G_{2})\geq 1,\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(N[x_{3}])-2\geq 0$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(x_{1},x_{2})-1\geq 1$. Because $S(G_{4})\geq 1$, these subproblems $G_{1},G_{2},G_{3},G_{4}$ have drops $(2.5,6),(3,7),(3.5,8),(2,5)$. We need:
@@
Line 7. After Line 1, subproblems $G-N[x_{1}]$ and $G-x_{1}-N[x_{2}]$ and $G-\{x_{1},x_{2}\}$ have drops $(2.5,6),(3,7),(1,2)$ directly, and $G-x_{1}-N[x_{2}]$ has drop $(3,8)$ after simplification. We need:
@@ 8.3 Line 5: branching on a 6-vertex with high-degree neighbors
First observe that $H$ cannot have any triangle $v_{1},v_{2},v_{3}$; for, in that case, we would have $S(v_{4},v_{5})\geq 1$ where $v_{4},v_{5}$ are any other 6-neighbors of $u$; we would branch on $v_{4},v_{5}$ at Line 3.
@@
If $H$ is a matching with exactly one edge $(v_{1},v_{2})$, with three other 6-neighbors $v_{3},v_{4},v_{5}$, we would have $S(v_{3},v_{4},v_{5})\geq 1$ due to the 3-triangle $u,v_{1},v_{2}$ and we would branch Line 4.
If $H$ is a matching with exactly two edges, suppose without loss of generality they are $(v_{1},v_{2}),(v_{3},v_{4})$ where $\deg(v_{4})=6$. If $\deg(v_{5})=\deg(v_{6})=6$, then again have $S(v_{4},v_{5},v_{6})\geq 1$ due to 3-triangle $u,v_{1},v_{2}$, and we would branch at Line 4. So suppose without loss of generality that $\deg(v_{5})=5$, in which case the ordering $v_{1},\dots,v_{6}$ satisfies Case II.
@@
Next suppose $H$ is not a matching, and some vertex $v_{4}\in H$ has neighbors $v_{5},v_{6}\in H$ where the other neighbors $v_{1},v_{2},v_{3}$ all have degree 6. Since $H$ has no triangle, we order these vertices so that $v_{2}\not\sim v_{3}$, satisfying Case III.
Finally, suppose $H$ is not a matching and some $v_{4}\in H$ has neighbors $v_{5},v_{6}\in H$ where $\deg(v_{4})=\deg(v_{5})=\deg(v_{6})=6$. The induced subgraph $H^{\prime}=G[\{v_{1},v_{2},v_{3}\}]$ must be a matching, as otherwise it would be covered by the argument in the preceding paragraph. So let $x$ be a vertex which is isolated in $H^{\prime}$. If $\deg(x)=6$, we set $v_{1}=x$ and $v_{2}$ to be the other 5-vertex (if any) and $v_{3}$ to be the 6-vertex. If $\deg(x)=5$, we set $v_{1},v_{2}$ to be the 6-neighbors of $u$ and $v_{3}=x$. ∎
@@
Case Ia: $\boldsymbol{G-\{v_{1},v_{2}\}-N[v_{3}]}$ is 3-vul. Split on $v_{1},\dots,v_{4}$, generating subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}],G_{4}=G-\{v_{1},v_{2},v_{3}\}-N[v_{4}],G_{5}=G-\{v_{1},v_{2},v_{3},v_{4}\}$. By Observation 62, we have $\operatorname{minsurp}(G_{1})\geq 2$, $\operatorname{minsurp}(G_{2})\geq 1$, $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(N[v_{3}])-2\geq 0$, $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(N[v_{4}])-3\geq-1$ and $\operatorname{minsurp}(G_{5})\geq\operatorname{shad}(v_{1},v_{2},v_{3})-1\geq 1$. So $G_{1},G_{2},G_{3},G_{4},G_{5}$ have drop $(2.5,6),(3,7),(3.5,8),(3.5,9),(2,4)$ directly. We can apply Proposition 31 to $G_{3}$, and have $S(G_{5})\geq 1$ due to 2-vertex $u$. We need:
@@
Case Ib: Split on $v_{1},v_{2},v_{3},v_{4}$, and apply (B-Pair) to vertex $u$ in the subproblem $G-\{v_{1},v_{2},v_{3},v_{4}\}$. This generates subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}],G_{4}=G-\{v_{1},v_{2},v_{3}\}-N[v_{4}],G_{0}=G-N[u]$ and $G_{5}=G-\{v_{1},v_{2},v_{3},v_{4}\}-N[v_{5},v_{6}]$.
Since we are after Line 1, subproblems $G_{0},G_{1},G_{2},G_{3}$ have drops $(2.5,6),(2.5,6),(3,7),(3.5,8)$ directly. We must have $\operatorname{minsurp}(G_{4})\geq 0$, as otherwise $\operatorname{shad}(v_{1},v_{2},N[v_{4}])\leq 0$, which would put us into Case Ia (with swapping the order of $v_{3}/v_{4}$). So $G_{4}$ has drop $(4,9)$.
For subproblem $G_{5}$, let $d=\deg_{G-N[v_{5}]}(v_{6})$; by Observation 62 we have $d\geq 3$. We claim that $\operatorname{minsurp}(G_{5})\geq 2-d$. For, consider an indset $I$ with $\operatorname{surp}_{G_{5}}(I)\leq 1-d\leq-2$; necessarily $|I|\geq 2$. Then $\operatorname{surp}_{G-\{v_{1},v_{2}\}-N[v_{5}]}(I\cup\{v_{6}\})\leq(1-d)+(d-1)+2=2$. So $G-\{v_{1},v_{2}\}-N[v_{5}]$ is 3-vul, putting us into Case Ia with alternate vertex ordering $v_{1}^{\prime}=v_{1},v_{2}^{\prime}=v_{2},v_{3}^{\prime}=v_{5}$. So $G_{5}$ is principal with excess two and $\Delta k=10+d\geq 13$ and $\Delta\mu\geq 5$.
@@
Case IIa: $\boldsymbol{G-v_{6}-N[v_{1}]}$ is 3-vul. Split on $v_{6},v_{1},v_{5}$, obtaining subproblems $G_{1}=G-N[v_{6}],G_{2}=G-v_{6}-N[v_{1}],G_{3}=G-\{v_{1},v_{6}\}-N[v_{5}],G_{4}=G-\{v_{1},v_{6},v_{5}\}$. By Observation 62, we have $\operatorname{minsurp}(G_{1})\geq 2,\operatorname{minsurp}(G_{2})\geq 1,\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(v_{6},N[v_{5}])-1\geq 0,\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(v_{1},v_{6})-1\geq 1$. So $G_{1},G_{2},G_{3},G_{4}$ have drops $(2.5,6),(3,7),(3,7),(1.5,3)$ directly. We apply Proposition 31 to $G_{3}$ and we have $S(G_{4})\geq 1$ (due to 3-triangle $u,v_{3},v_{4}$), so overall we need:
@@
Case IIb: Split on $v_{6},v_{4},v_{3}$ and apply (B-Fun2) to vertex $v_{1}$ in subproblem $G-\{v_{6},v_{4},v_{3}\}$. Overall, this generates subproblems $G_{1}=G-N[v_{6}],G_{2}=G-v_{6}-N[v_{4}],G_{3}=G-v_{6}-N[v_{3}],G_{4}=G-\{v_{3},v_{4},v_{6},v_{1}\}$ and $G_{5}=G-\{v_{3},v_{4},v_{6}\}-N[v_{1},v_{5}]$. By Observation 62, we have $\operatorname{minsurp}(G_{1})\geq 2$ and $\operatorname{minsurp}(G_{2})\geq 1$ and $\operatorname{minsurp}(G_{3})\geq 1$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(v_{4},v_{6})-2\geq 0$. So the subproblems have drops $(2.5,6),(3,7),(3,7),(2,4)$ directly; subproblem $G_{4}$ has a 2-vertex $u$, so it has drop $(2,5)$ after simplification.
For subproblem $G_{5}$, let $d=\deg_{G-v_{6}-N[v_{1}]}(v_{5})$. We have $d\geq 3$, as otherwise $S(v_{6},N[v_{1}])\geq 1$ and it would be covered in Case IIa. We must have $\operatorname{minsurp}(G_{5})\geq 2-d$; for, if we have an indset $I$ with $\operatorname{surp}_{G_{5}}(I)\leq 1-d\leq-2$, where necessarily $|I|\geq 2$, then $\operatorname{surp}_{G-v_{6}-N[v_{1}]}(I\cup\{v_{5}\})\leq 2$ and $G-v_{6}-N[v_{1}]$ is 3-vul; it would be covered in Case IIa.
@@
Case III. Split on $v_{1},v_{2},v_{3}$, obtaining subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}],G_{4}=G-\{v_{1},v_{2},v_{3}\}$. By Observation 62, we have $\operatorname{minsurp}(G_{1})\geq 2$ and $\operatorname{minsurp}(G_{2})\geq 1$ and $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(v_{1},N[v_{3}])-1\geq 0$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(v_{1},v_{2})-1\geq 1$. So the subproblems have drops $(2.5,6),(2.5,6),(3,7),(1.5,3)$ directly; furthermore, $G_{4}$ has a kite $u,v_{4},v_{5},v_{6}$ and so it has net drop $(2,5)$ by Lemma 14. We need: